# Puffy Venuses: the Mass - Radius Impact of Carbon-Rich Atmospheres on Lava Worlds

### v1

@@ 3 Results
We report how the three observables of puffy Venuses, i.e., their mass, radius (or equivalently, assuming an Earth-like interior, the atmospheric height), and the gas composition at the top of their atmospheres, are impacted by the equilibrium temperature, the mantle redox state, and the bulk carbon content. We also investigate the effect of different thermochemical scenarios on the mass-radius (M-R) relations. For our fiducial case (fig. 1a), we model planets with an $T_{eq}$ = 1500 K, and fix their bulk volatile C and H content to that of the ordinary chondrites (OC), with 1200 ppm C and 459 ppm H (Schaefer & Fegley, 2017).
@@ 3.1 Mass - Radius relations
We calculate the radii of puffy Venuses with masses between 0.5 - 15 $M_{\oplus}$ for our fiducial case, (Fig. 2), and subsequently probe different equilibrium temperatures: $T_{eq}=$[1000, 1500, 2000] K (fig. 2a).
We find that, as expected, radius inflation is the most pronounced for highly-irradiated, low-mass planets. For instance, an Earth-mass planet can host an atmosphere that is 0.3 $R_{\oplus}$ high, with $T_{eq}$ of 2000 K (fig. 2a). This is a few times larger than the uncertainties in the best mass-radius observations. Therefore, these puffy Venus planets follow a distinct M-R relation to their terrestrial counterparts. An alternative explanation for underdense rocky worlds is that they form iron-depleted (Plotnykov & Valencia, 2020). We compare the excess radius caused by this scenario (fig 2a, shaded region) with the M-R relations we calculate, and show that the puffy Venus scenario can produce greater radius inflation than Fe-poor rocky worlds for low-mass and irradiated planets. Specifically, a carbon-rich atmosphere can produce greater puffiness in $<4M_{\oplus}$ planets with $T_{eq}$ = 2000 K, or $<2M_{\oplus}$ planets at 1500 K. Otherwise these two scenarios are degenerate.
One can compare these M-R relations to exoplanets of similar $T_{eq}$. For planets that are below the M-R, and have similar $T_{eq}$ or hotter, a carbon-rich atmosphere can explain their excess radius. We identify 9 puffy Venus candidates: GJ 1252b, TOI-500 b, Kepler-10 b, Kepler-36 b, HD 219134 b, HD 3167 b, WASP-47 e, 55 Cnc e, and TOI-561 b. The last two are of particular interest. TOI-561 b’s combination of high irradiation and low-density disfavors other interior scenarios, such as having H2 or H2O layers, or a lower Fe/Si ratio than its host star. While recent JWST observation on the eclipse of 55 Cnc e indicates active heat redistribution, favoring a CO-CO2 atmosphere (Hu et al. 2024, see Section 4).
@@
We quantify the effect of different bulk carbon and hydrogen abundances on the M-R relations (fig. 2b) using volatile C and H abundances characteristic of Solar-System materials: bulk Earth, ordinary chondrites, enstatite chondrites, and CI carbonaceous chondrites, see Table 1. For each volatile C abundance, we contrast the less-insolated $T_{eq}=$ 1000 K case with a hotter $T_{eq}=$ 2000 K case. For the 1000 K case, we do not display the M-R of wet CI bulk abundances. This is because the wet CI abundances generate high MO surface pressure, which leads to graphite saturation in $>2\mathrm{M_{\oplus}}$ planets (see B.2), for which our model does not apply. We demonstrate that higher volatile C and H abundances leads to greater atmospheric height, especially for low-mass planets $\lesssim 2M_{\oplus}$. However, the temperature effect on puffiness is stronger - only the most carbon-rich composition for the $T_{eq}$ = 1000 K case can reach similar puffiness as its hotter counterparts at $\geq 2M_{\oplus}$.
We test for the effect of chemical quenching and neglecting atmospheric thermochemistry on the M-R relations by comparing scenarios I, II and III in fig. 2c. In general, assuming pure CO2-H2O atmospheres (III) lead to the least radius inflation while adopting a uniform atmosphere chemically equilibrated at the MO surface (II), predicts the most extensive atmospheres. Our fiducial scenario I, which enforces chemical equilibrium locally throughout the atmosphere, predicts an M-R relation somewhere in between those predicted by the previous two. Yet again, the variation in M-R relations due to $T_{eq}$ (orange vs. blue lines) is more substantial than the M-R variations from different atmospheric chemistry scenarios.
@@
Further, if we assume the atmosphere is convective down to the MO surface rather than allowing for a deep radiative zone (fig. 1d, orange solid vs dashed), then $T_{s}$ is drastically overestimated by a factor of 2. This increases $z_{atm}$ by 37%. A fully convective atmosphere can be interpreted as (i) the atmosphere having a low $\gamma$, the opacity ratio of incoming, stellar radiation and outgoing, thermal radiation (Guillot, 2010), or ii the planet having a high internal heat flux. This is relevant to atmospheres hosting extremely efficient greenhouse gases, or young planets still hot from formation. Our choices of parameters, in contrast, assume a maximum $\gamma$ and no internal heating (see Appendix B.1), representing an evolved world. Thus the realistic $z_{atm}$ of a puffy Venus is likely bracketed by our fiducial value and that of the fully convective case (fig. 2a, solid vs. dashed orange lines), assuming our choice of thermal opacity $\kappa_{th}$ is realistic. As we further discuss in Appendix B.1, our conclusions are qualitatively robust to the exact choices of opacities.
Chemically active atmospheres are thicker than pure CO2 - H2O ones because CO2 thermally decomposes into CO and O2 at $\gtrsim$ 4000 K in the convective layer. CO-rich atmospheres are then puffier for two reasons. Firstly, CO is about $\sim$2 orders of magnitude less soluble than CO2 at P $\lesssim$ 1 kBar. Thermal dissociation of CO2 into CO and O2 at high temperatures thus favors carbon partitioning into the atmosphere. Furthermore, CO2, as a larger molecule, has a higher specific heat capacity than CO, resulting in a shallower adiabatic lapse rate (eqn. A34). This means that for the same tropopause P-T conditions, a CO-dominated atmosphere would have a steeper temperature gradient and reach higher $T_{s}$ for similar MO surface pressures (see, for instance, fig. 1d, planet a vs. b). These factors facilitate a higher MO surface temperature - and thus a more extended atmosphere - if an atmosphere receives enough insolation to achieve some CO2 dissociation at its base (see planet c, fig. 1 for example). Parallel to CO2 and CO, H2O also dissociates at high temperature to H2 and O2, inflating the atmosphere (see e.g. fig. 1c). However, the H species’ contribution is small at our fiducial C and H abundances and atmospheric C/O since they only account for a few percent of the atmosphere by mole. Holding the C/O constant but adding more H would make the atmosphere (chemically) reduced, resulting in a hydrocarbon-CO-H2 atmosphere, which resembles a primary or hybrid atmosphere (Tian & Heng, 2024). Finally, we note that the deep, isothermal layer sees some reversion of the CO2 and H2O’s thermal decomposition, but this is not enough to neutralize the trends in the convective layer.
@@ 3.2 Mantle redox effects on volatile speciation and atmospheric height
The redox states of exoplanet MOs are not well constrained and might be influenced by internal processes such as core formation (Wade & Wood, 2005; Lichtenberg, 2021) and MO (partial) solidification (Hirschmann, 2012, 2022; Maurice et al., 2023). Yet the MO’s redox state strongly impacts the height and composition of a puffy Venus’ atmosphere. We probe this effect by varying the atmospheric C/O ratio of our fiducial model between 0.995 and 0.3, corresponding to moderately reduced to highly oxidized MO’s (fig. 3). As shown in fig. 3b, varying atmospheric C/O ratio significantly impacts the atmospheric height, $z_{atm}$, by up to a factor of $\sim$3. $z_{atm}$ reaches a minimum at C/O = 0.5, and a maximum at the reduced end where C/O = 0.995. Surface temperature $T_{s}$ follows a similar trend, varying from $\sim$4400 K at C/O =0.5 to $\sim$5600 K near the reduced end. Thus, the radius inflation we found assuming solar C/O = 0.55 in fig. 2 are conservative estimates.
The atmospheric composition (fig. 3a) changes as a direct consequence of varying atmospheric C/O. At the MO surface, CO is the dominant species for a wide span of reduced atmospheres (C/O $\sim$ 0.6 - 1), while only at the vicinity of C/O $\sim$ 0.5 does CO2 become dominant. At the oxidized extreme (C/O $\lesssim$ 0.32), the atmosphere becomes O2 dominated, while H2 only plays a significant role at the most reduced end. In terms of the volatile speciation between the MO and the atmosphere (fig. 3c, d), the vast majority of H is in the form of H2O dissolved in the MO, while the majority of C remains in the atmosphere, as dictated by the solubility relations. Therefore, although the bulk H reservoir is larger (H/C = 4.59 by moles), H contributes little to the atmosphere, except at the most reducing end. The deep atmosphere composition (fig. 3a) follows overall similar trends as the top-of-atmosphere composition (fig. 3d), with the important difference that CO2 is overall less prominent in the deep atmosphere, while O2 and CO are more present. This is the result of thermal decomposition.
@@
Finally, although in the context of the Solar System’s chondritic record, OC has one of the lowest volatile contents, our fiducial puffy Venuses of OC-like abundances could still host a massive, high molecular mass atmosphere with extreme surface conditions - total pressure up to $\sim$0.6 GPa and temperatures of $\gtrsim$6000 K. We discuss the caveats for applying our model to these extreme P-T regimes in Appendix B.4.
@@ 4 Discussion
Our results establish that puffy Venuses constitute an alternative interpretation for low-density terrestrial exoplanets. This interpretation of rocky exoplanet interiors offers some unique advantages. We illustrate the arguments for a puffy Venus interpretation with the cases of TOI-561 b and 55 Cancri e.
@@ 4.1 The puffy Venus case for TOI-561 b
The super-Earth TOI-561 b has garnered recent observational interest (Weiss et al., 2021; Lacedelli et al., 2021, 2022; Brinkman et al., 2023; Patel et al., 2023). As an ultra-short-period planet with $T_{eq}\geq 2000$ K, TOI-561 b has one of the lowest densities among its peers ($\rho_{b}=4.3\pm 0.5$ g/cc, Patel et al. (2023)). While Weiss et al. (2021) suggest the planet can be consistent with both a metal-poor, mantle-only interior or an Earth-like one, follow-up observations have established its bulk density too low to be Earth-like (Brinkman et al., 2023; Patel et al., 2023), favoring the mantle-only scenario (Fig. 2a, rocky threshold radius line, RTR). Lacedelli et al. (2021, 2022) on the other hand, suggests incorporating $>$ 30 wtof water ice, necessitating a formation scenario with migration from beyond the snow line to its current location. Brinkman et al. (2023) also suggests the possibility of a high mean-molecular-weight (MMW) atmosphere, such as that of water vapor, CO2, or rock vapor. Finally (Patel et al., 2023) reported a TESS secondary eclipse depth of $27.4\pm 11$ ppm, consistent with a rock vapor atmosphere, although the accuracy is not good enough for further inferences.
@@
Some of these interior scenarios have their drawbacks. Although the host star, TOI-561, has a somewhat depleted Fe/Si ratio compared to the Sun (0.60 vs 0.89 by moles, Weiss et al. (2021)), suggesting the planet’s solid portion could have a lower bulk density. Yet using the M-R relations from Plotnykov & Valencia (2020), we estimated that this effect could only account for $\sim$ 20% of TOI-561 b’s excess radius. Therefore, if TOI-561 were a naked-rock planet, it would still require an abnormally metal-depleted mixture. Such a scenario is challenging to account for from a formational standpoint (Scora et al., 2020, 2022). Whereas a water-rich envelope at equilibrium temperatures of 2000 K would require very little water (Turbet et al., 2019; Dorn & Lichtenberg, 2021; Pierrehumbert, 2023), and given the fact that this type of atmosphere would be highly susceptible to XUV-driven photolysis and escape (Tian, 2009, 2015; Wordsworth & Pierrehumbert, 2013; Bolmont et al., 2017), the water most likely need to be constantly replenished.
@@
Instead, a puffy Venus interpretation for TOI-561 b may be more straightforward to justify. As our M-R relations (fig. 2) indicate, a bulk volatile carbon content available to dry, inner Solar System chondrite can produce more significant radius inflation than the planet requires, assuming an Earth-like interior - no formation beyond the snow line necessary. Further, since the carbonic gases are less soluble than water, a puffy Venus TOI-561 b requires less bulk volatile content - ($<$ 1 wt%) than a steam atmosphere one ($\sim$ 5 wt%, Dorn & Lichtenberg (2021)), and is less susceptible to atmospheric loss. Our results (fig. 3) show that, even when the hydrogen budget is somewhat greater by moles (H/C = 4.59 for ordinary chondrite), the atmosphere stays carbon-dominated for a wide range of mantle redox states.
@@
One of the most observed super-Earths, 55 Cancri e, has a bulk density marginally consistent with a purely rocky interior (fig. 4). Pre-JWST observations of its phase curve and transmission spectroscopy had been inconclusive in detecting an atmosphere (see reviews in Bourrier et al. 2018 and Demory et al. 2023). Neither a low mean-molecular-weight atmosphere (Esteves et al., 2017; Jindal et al., 2020; Deibert et al., 2021; Zhang et al., 2021), nor an extended rock vapor atmosphere (Tabernero et al., 2020; Keles et al., 2022) is favored by observation.
@@
Analogous to those of TOI-561 b, previous interior models of 55 Cancri e favor some combination of a metal-depleted interior, a modest ice layer, and a small atmosphere of up to $\sim$8its radius (Dorn et al., 2017; Crida et al., 2018; Bourrier et al., 2018). These degenerate scenarios, together with the poorly-constrained stellar refractory element ratios, result in inconclusive inferences of the atmosphere’s metallicity and the mantle composition (Crida et al., 2018; Bourrier et al., 2018). More exotic interior scenarios include a carbide-rich mantle motivated by 55 Cnc’s super-solar C/O (Madhusudhan et al., 2012). Heng (2023) envisions a tenuous, transient, outgassed atmosphere, which seeks to explain the variability in 55 Cnc e’s secondary eclipse depths.
@@
Recently, Hu et al. (2024) obtained a thermal emission spectrum of 55 Cnc e from JWST, which indicates a volatile atmosphere rich in CO2 or CO. This discovery strongly promotes the puffy Venus scenario for this planet. In fig. 4, we probe four possible C, H abundances for 55 Cnc e, ranging from inner-Solar System enstatite chondrite (EH) to the dry CI carbonaceous chondrite. We also test solar C/O = 0.55 and reduced C/O = 0.78 (solid vs dashed lines, fig. 4). This range is motivated by the different measurements on the host star’s C/O (Teske et al., 2013; Brewer & Fischer, 2016).

### v2

@@ 3 Results
Figure 1: Three sample atmospheric profiles of Earth-mass planets with OC total C and H abundances. a (fiducial): Atmospheric composition profile of a planet with solar C/O = 0.55 and $T_{eq}$ = 1500 K. The shaded region denotes the convective layer. b: Same as a but with a more reducing atmosphere of C/O = 0.8. c: Same as a but with $T_{eq}=2000$ K. d: The atmosphere pressure-temperature of planets a (solid green), b (dash-dot green) and c (solid orange). For comparison, the dashed orange line is an atmosphere at $T_{eq}=2000$ K that stays convective to the MO surface. Both a higher $T_{eq}$ and a more reduced environment promote higher surface temperatures, resulting in larger atmospheres. For reference, $z_{atm}$= 0.162, 0.233, 0.307 $R_{\oplus}$ for planets a, b and c, respectively.
Figure 2: Mass - radius relations of puffy Venuses. a: M-R relations compared to observed exoplanets. The grey line is the M-R for terrestrial planets with an Earth-like core mass fraction of 0.325 (Plotnykov & Valencia, 2020). Colored solid lines are puffy Venus radii assuming different $T_{eq}$, with the fiducial, OC-like volatile C and H abundances. The orange dashed line is the M-R for $T_{eq}=2000$ K planets with a convective deep atmosphere to the MO surface instead of a deep radiative zone. The shaded region is the range of M-R space an air-less rocky world can occupy for all possible Fe/Si $\in[0,1]$. Planets less dense than this region require a volatile component. Measured exoplanet samples are from the NASA Exoplanet Archive (2023) and only contain those with relative errors below 30%. The nine planets highlighted with black circles are our preferred puffy Venus candidates. In rising mass order, they are GJ 1252 b, TOI-500 b, TOI-561 b, Kepler-10 b, Kepler-36 b, HD 219134 b, HD 3167 b, WASP-47 e, and 55 Cnc e. TOI-561 b’s radius is from Patel et al. (2023); its mass is from Brinkman et al. (2023) b: M-R of puffy Venuses with a range of chondritic C abundances, more opaque lines correspond to less C. They are color-coded by temperature - $T_{eq}$ = 1000 K (blue) and 2000 K (orange). The dashed orange line is for wet CI composition, while the solid line with the same color is for dry CI composition. The dotted portion of dry CI, 1000 K line are atmospheres that should experience graphite precipitation; see text. c: M-R of models assuming different chemical scenarios - I (solid), local chemical equilibrium throughout the atmosphere, II (dash-dot), chemical equilibrium at the MO surface, and a CO2 + H2O atmosphere (dotted). Fiducial OC C and H abundances are used. Same color-coding as b. The solid 1000 K and 2000 K lines in a, the 2 OC lines in b, and the 2 Scenario I lines in c are the same.
We report how the three observables of puffy Venuses, i.e., their mass, radius (or equivalently, assuming an Earth-like interior, the atmospheric height), and the gas composition at the top of their atmospheres are impacted by the equilibrium temperature, the mantle redox state, and the bulk carbon content. We also investigate the mass - radius effect of different thermochemical scenarios: (I) local chemical equilibrium, (II) quenched at MO surface and (III) CO2 + H2O atmospheres (see Section 2). For our fiducial case (fig. 1a), we model planets with an $T_{eq}$ = 1500 K and fix their bulk volatile C and H content to that of the ordinary chondrites (OC), with 1200 ppm C and 459 ppm H (Schaefer & Fegley, 2017).
@@ 3.1 Mass - Radius relations
We calculate the radii of puffy Venuses with masses between 0.5 - 15 $M_{\oplus}$ for our fiducial case, Fig. 2, and probe different equilibrium temperatures: $T_{eq}=$ [1000, 1500, 2000] K (fig. 2a). Our key result is the significant difference of puffy Venus M-R relations with that of Earth-like planets (Fig. 2a, colored lines vs. grey line).
We find that, as expected, radius inflation is the most pronounced for highly-irradiated, low-mass planets. For instance, an Earth-mass planet can host an atmosphere that is 0.3 $R_{\oplus}$ high, with $T_{eq}$ of 2000 K (fig. 2a). This is a few times larger than the uncertainties in the best mass-radius observations. Therefore, these puffy Venus planets follow a distinct M-R relation to their terrestrial counterparts. An alternative explanation for underdense rocky worlds is that they form iron-depleted (Plotnykov & Valencia, 2020). We compare the excess radius caused by this Scenario (fig 2a, shaded region) with the M-R relations we calculate, and show that the puffy Venus scenario can produce greater radius inflation than Fe-poor rocky worlds for low-mass and irradiated planets. Specifically, a carbon-rich atmosphere can produce greater puffiness in $<4M_{\oplus}$ planets with $T_{eq}$ = 2000 K, or $<2M_{\oplus}$ planets at 1500 K. Otherwise these two scenarios are degenerate in the mass-radius space, although atmospheric characterization can help break this degeneracy.
To identify which exoplanets can be puffy Venuses, we need to compare their data to these M-R relations at similar $T_{eq}$. For planets that are below the M-R and have similar or higher $T_{eq}$, a carbon-rich atmosphere can explain their excess radius. We identify 9 puffy Venus candidates: GJ 1252b, TOI-500 b, Kepler-10 b, Kepler-36 b, HD 219134 b, HD 3167 b, WASP-47 e, 55 Cnc e, and TOI-561 b. The last two are of particular interest. TOI-561 b’s combination of high irradiation and low-density disfavors other interior scenarios, such as having H2 or H2O layers, or a lower Fe/Si ratio than its host star. While recent JWST observations on the eclipse of 55 Cnc e indicate active heat redistribution and favored a CO-CO2 atmosphere (Hu et al. 2024, see Section 4).
@@
Having a baseline for the structural effects of carbon-rich hot atmospheres, we quantify the effect of different bulk carbon and hydrogen abundances on the M-R relations (fig. 2b) using volatile C and H abundances characteristic of Solar-System materials: bulk Earth, ordinary chondrites, enstatite chondrites, and CI carbonaceous chondrites, see Table 1. For each volatile C abundance, we contrast the less-insolated $T_{eq}=$ 1000 K case with a hotter $T_{eq}=$ 2000 K case. For the 1000 K case, we do not display the M-R of wet CI bulk abundances. This is because the wet CI abundances generate high MO surface pressure, which leads to graphite saturation in $>2\mathrm{M_{\oplus}}$ planets (see 4.3.2), for which our model does not apply. We demonstrate that higher volatile C and H abundances lead to greater atmospheric height, especially for low-mass planets $\lesssim 2M_{\oplus}$. However, the temperature effect on radius is stronger – only the most carbon-rich composition for the $T_{eq}$ = 1000 K case can reach similar puffiness as its hotter counterparts at $\geq 2M_{\oplus}$.
We test for the effect of (a) assuming a vertically uniform atmosphere composition (i.e., chemical quenching) and (b) neglecting atmospheric thermochemistry on the M-R relations by comparing scenarios I, II, and III in fig. 2c. In general, assuming pure CO2-H2O atmospheres (III) lead to the least radius inflation, while adopting a uniform atmosphere chemically equilibrated at the MO surface (II) predicts the most extensive atmospheres. Our fiducial Scenario I, which enforces chemical equilibrium locally throughout the atmosphere, predicts an M-R relation somewhere in between those predicted by the other two. Yet again, the variation in M-R relations due to $T_{eq}$ (orange vs. blue lines in Fig. 2c) is more substantial than the M-R variations from different atmospheric chemistry scenarios.
@@
Further, if we assume the atmosphere is convective down to the MO surface rather than allowing for a deep radiative zone (fig. 1d, orange solid vs dashed), then $T_{s}$ is drastically overestimated by a factor of 2. This increases $z_{atm}$ by 37%. A fully convective atmosphere can be interpreted as (i) the atmosphere having a low $\gamma$, the opacity ratio of incoming stellar radiation and outgoing thermal radiation (Guillot, 2010), or (ii) the planet having a high internal heat flux. This is relevant to atmospheres hosting extremely efficient greenhouse gases or young planets still hot from formation. Our choices of parameters, in contrast, assume a maximum $\gamma$ and no internal heating (see Section 4.3.1), representing an evolved world. Thus, the realistic $z_{atm}$ of a puffy Venus is likely bracketed by our fiducial value and that of the fully convective case (fig. 2a, solid vs. dashed orange lines), assuming our choice of thermal opacity $\kappa_{th}$ is realistic. As we further discuss in Section 4.3.1, our conclusions are qualitatively robust to the exact choices of opacities.
Chemically active atmospheres are thicker than pure CO2 - H2O ones because CO2 thermally decomposes into CO and O2 at $\gtrsim$ 4000 K in the convective layer. CO-rich atmospheres are then puffier for two reasons. Firstly, CO is about $\sim$2 orders of magnitude less soluble than CO2 at P $\lesssim$ 1 kBar. Thermal dissociation of CO2 into CO and O2 at high temperatures thus favors carbon partitioning into the atmosphere. Furthermore, CO2, as a larger molecule, has a higher specific heat capacity than CO, resulting in a shallower adiabatic lapse rate (Equation 33). This means that for the same tropopause P-T conditions, a CO-dominated atmosphere would have a steeper temperature gradient and reach higher $T_{s}$ for similar MO surface pressures (see, for instance, fig. 1d, planet a vs. b). These factors facilitate a higher MO surface temperature - and thus a more extended atmosphere - if an atmosphere receives enough insolation to achieve some CO2 dissociation at its base (see planet c, fig. 1 for example). Parallel to CO2 and CO, H2O also dissociates at high temperatures to H2 and O2, inflating the atmosphere (see, e.g., fig. 1c). However, the H species’ contribution is small at our fiducial C and H abundances and atmospheric C/O since they only account for a few percent of the atmosphere by mole. Holding the C/O constant but adding more H would make the atmosphere chemically reduced, resulting in a hydrocarbon-CO-H2 atmosphere, which resembles a primary or hybrid atmosphere (Tian & Heng, 2024). Finally, we note that the deep, isothermal layer sees some reversion of the CO2 and H2O’s thermal decomposition, but this is not enough to neutralize the trends in the convective layer.
@@ 3.2 Mantle redox effects on volatile speciation and atmospheric height
The redox states of exoplanet MOs are not well constrained and might be influenced by internal processes such as core formation (Wade & Wood, 2005; Lichtenberg, 2021) and MO (partial) solidification (Hirschmann, 2012, 2022; Maurice et al., 2023). Yet the MO’s redox state strongly impacts the height and composition of a puffy Venus’ atmosphere. We probe this effect by varying the atmospheric C/O ratio of our fiducial model between 0.995 and 0.3, corresponding to moderately reduced to highly oxidized MO’s (fig. 3). As shown in fig. 3b, varying atmospheric C/O ratio significantly impacts the atmospheric height, $z_{atm}$, by up to a factor of $\sim$3. $z_{atm}$ reaches a minimum at C/O = 0.5, and a maximum at the reduced end, where C/O = 0.995. Surface temperature $T_{s}$ follows a similar trend, varying from $\sim$4400 K at C/O =0.5 to $\sim$5600 K near the reduced end. Thus, the radius inflation we found assuming solar C/O = 0.55 in fig. 2 are conservative estimates.
The atmospheric composition changes as a direct consequence of varying atmospheric C/O (see fig. 3a). At the MO surface, CO is the dominant species for a wide span of reduced atmospheres (C/O $\sim$ 0.6 - 1), while only at the vicinity of C/O $\sim$ 0.5 does CO2 become dominant. At the oxidized extreme (C/O $\lesssim$ 0.32), the atmosphere becomes O2 dominated, while H2 only plays a significant role at the most reduced end. In terms of the volatile speciation between the MO and the atmosphere (fig. 3c, d), the vast majority of H is in the form of H2O dissolved in the MO, while the majority of C remains in the atmosphere, as dictated by the solubility relations. Therefore, although the bulk H reservoir is larger (H/C = 4.59 by moles), H contributes little to the atmosphere, except at the most reducing end. The deep atmosphere composition (fig. 3a) follows overall similar trends as the top-of-atmosphere composition (fig. 3d), with the important difference that CO2 is overall less prominent in the deep atmosphere, while O2 and CO are more present. This is the result of thermal decomposition.
@@
Finally, although in the context of the Solar System’s chondritic record, OC has one of the lowest volatile contents, our fiducial puffy Venuses of OC-like abundances could still host a massive, high molecular mass atmosphere with extreme surface conditions - total pressure up to $\sim$0.6 GPa and temperatures of $\gtrsim$6000 K. We discuss the caveats for applying our model to these extreme P-T regimes in Section 4.3.4.
@@ 4 Discussion
Our results suggest that puffy Venuses constitute an alternative interpretation for underdense terrestrial exoplanets. This interpretation of rocky exoplanet interiors offers some unique advantages. We illustrate the arguments for a puffy Venus interpretation with the cases of TOI-561 b and 55 Cancri e.
@@ 4.1 The puffy Venus case for TOI-561 b
The super-Earth TOI-561 b has garnered recent observational interest (Weiss et al., 2021; Lacedelli et al., 2021, 2022; Brinkman et al., 2023; Patel et al., 2023). As an ultra-short-period planet with $T_{eq}\geq 2000$ K, TOI-561 b has one of the lowest densities among its peers ($\rho_{b}=4.3\pm 0.5$ g/cc, Patel et al. (2023)). While Weiss et al. (2021) suggested the planet can be consistent with both a iron-poor, mantle-only interior or an Earth-like one, follow-up observations have established its bulk density too low to be Earth-like (Brinkman et al., 2023; Patel et al., 2023), favoring the mantle-only Scenario (Fig. 2a, rocky threshold radius line, RTR). Lacedelli et al. (2021, 2022) on the other hand, suggested incorporating $>$ 30 wtof water ice, necessitating a formation scenario with migration from beyond the snow line to its current location. Brinkman et al. (2023) also qualitatively argued for the possibility of a high mean-molecular-weight (MMW) atmosphere, such as that of water vapor, CO2, or rock vapor. Finally, (Patel et al., 2023) reported a TESS secondary eclipse depth of $27.4\pm 11$ ppm, consistent with a rock vapor atmosphere, although the accuracy is not good enough for further inferences.
@@
Some of these interior scenarios have their drawbacks. Although the host star, TOI-561, has a somewhat depleted Fe/Si ratio compared to the Sun (0.60 vs 0.89 by moles, Weiss et al. 2021), suggesting the planet’s solid portion could have a lower bulk density. Yet using the M-R relations from Plotnykov & Valencia (2020), we estimate that this effect could only account for $\sim$ 20% of TOI-561 b’s excess radius. Therefore, if TOI-561 were a naked-rock planet, it would still require an abnormally iron-depleted mixture. Such a scenario is challenging to account for from a formational standpoint (Scora et al., 2020, 2022). On the other hand, a water-rich envelope at equilibrium temperatures of 2000 K would require very little water (Turbet et al., 2019; Dorn & Lichtenberg, 2021; Pierrehumbert, 2023), and given the fact that this type of atmosphere would be highly susceptible to XUV-driven photolysis and escape (Tian, 2009, 2015; Wordsworth & Pierrehumbert, 2013; Bolmont et al., 2017), the water most likely needs to be constantly replenished.
@@
Instead, a puffy Venus interpretation for TOI-561 b may be more straightforward to justify. As our M-R relations (fig. 2) indicate, a bulk volatile carbon content available to dry, inner Solar System chondrite can produce more significant radius inflation than the planet requires, assuming an Earth-like interior - no formation beyond the snow line necessary. Further, since the carbonic gases are less soluble than water, a puffy Venus TOI-561 b requires less bulk volatile content - ($<$ 1 wt%) than a steam atmosphere one ($\sim$ 5 wt%, Dorn & Lichtenberg (2021)), and is less susceptible to atmospheric loss. Our results (fig. 3) show that, even when the hydrogen budget is somewhat greater by moles (H/C = 4.59 for ordinary chondrite), the atmosphere stays carbon-dominated for a wide range of mantle redox states. This qualitatively agrees with Sossi et al. (2023), who experimentally updated the solubility of water and modeled C-O-H atmospheres in equilibrium with MO’s but with lower bulk abundances than ours. Although their water solubility is up to $\sim$50lower than the one we adopt, they nevertheless found C-dominated atmospheres to be more common than steam-dominated ones.
@@
One of the most observed super-Earths, 55 Cancri e, has a bulk density marginally consistent with a purely rocky interior (fig. 4). Pre-JWST observations of its phase curve and transmission spectroscopy had been inconclusive in detecting an atmosphere (see reviews in Bourrier et al. 2018 and Demory et al. 2023). Neither a low mean-molecular-weight atmosphere (Esteves et al., 2017; Jindal et al., 2020; Deibert et al., 2021; Zhang et al., 2021), nor an extended rock vapor atmosphere (Tabernero et al., 2020; Keles et al., 2022) was favored by observation.
@@
Analogous to those of TOI-561 b, previous interior models of 55 Cancri e favored some combination of an iron-depleted interior, a modest ice layer, and a small atmosphere of up to $\sim$8its radius (Dorn et al., 2017; Crida et al., 2018; Bourrier et al., 2018). These degenerate scenarios, together with the poorly-constrained stellar refractory element ratios, resulted in inconclusive inferences of the atmosphere’s metallicity and the mantle composition (Crida et al., 2018; Bourrier et al., 2018). More exotic interior scenarios include a carbide-rich mantle motivated by 55 Cnc’s super-solar C/O (Madhusudhan et al., 2012). Heng (2023) envisioned a tenuous, transient, outgassed atmosphere, which seeks to explain the variability in 55 Cnc e’s secondary eclipse depths.
@@
Recently, Hu et al. (2024) obtained a thermal emission spectrum of 55 Cnc e from JWST, which indicated a volatile atmosphere rich in CO2 or CO. This discovery strongly promotes the puffy Venus scenario for this planet. In fig. 4, we probe four possible C, H abundances for 55 Cnc e, ranging from inner-Solar System enstatite chondrite (EC) to the dry CI carbonaceous chondrite. We also test solar C/O = 0.55 and reduced C/O = 0.78 (solid vs dashed lines, fig. 4). This range is motivated by the different measurements on the host star’s C/O (Teske et al., 2013; Brewer & Fischer, 2016).